OpenStudy (anonymous):

perimeter of rectangle = 34 length of diagonal = 13 width = x derive x^2 - 17x + 60 = 0 find the dimensions of the rectangle

OpenStudy (amistre64):

the optimal measurements?

OpenStudy (nikvist):

dimensions of rectangle are 5 and 12

OpenStudy (anonymous):

how did you get that? :(

OpenStudy (amistre64):

iwas just about to say 5 and 12.... lol

OpenStudy (anonymous):

you're both gods to me haha

OpenStudy (amistre64):

the 13 diagonal means that there is only a few setups of legs to match.... it just so happens that a 5-12-13 triangle works

OpenStudy (anonymous):

so why is there a quadratic equation in the problem?

OpenStudy (amistre64):

clutter probably :) maybe its part of the next problem

OpenStudy (nikvist):

\[2(a+b)=34\quad,\quad\sqrt{a^2+b^2}=13\] solve this system of equations

OpenStudy (amistre64):

you tend to use the derivative stuff to find the max or min area of the box..... but the 13 and 34 tend to limit you already

OpenStudy (anonymous):

thanks to both of you, but does the quadratic equation have anything to do with the problem

OpenStudy (amistre64):

it can.... if you dont know the limitations of the other stuff, that can help you to find a length...maybe

OpenStudy (anonymous):

thank you again, and nikvist too

OpenStudy (radar):

Here is how the \[x ^{2}-17x+60\] Came into play. Let x = dimension L The perimeter is 34. 34-2x = 2y which is the other dimension doubled 2y. Divide that by 2 getting: 17-x= width. Now use that to get the diagonal.

OpenStudy (anonymous):

The quadratic equation is there because you need to find the value of the width using the quadratic equation.

OpenStudy (radar):

\[x ^{2}+(17-x)^{2}=13^{2}=169\]

OpenStudy (radar):

\[x ^{2}+289-34x+x ^{2}=169\] \[2x ^{2}-34x+289=169\] \[2x ^{2}-34x+120=0\] \[x ^{2}-17x+60=0\] Did you follow

OpenStudy (radar):

You can factor solve it getting (x-5)(x-12)=0 x=5 and 12